### v1

# Subobject-level Image Tokenization

Abstract
Transformer-based vision models typically tokenize images into fixed-size square patches as input units, which lacks the adaptability to image content and overlooks the inherent pixel grouping structure. Inspired by the subword tokenization widely adopted in language models, we propose an image tokenizer at a subobject level, where the subobjects are represented by semantically meaningful image segments obtained by segmentation models (e.g., segment anything models). To implement a learning system based on subobject tokenization, we first introduced a Sequence-to-sequence AutoEncoder (SeqAE) to compress subobject segments of varying sizes and shapes into compact embedding vectors, then fed the subobject embeddings into a large language model for vision language learning. Empirical results demonstrated that our subobject-level tokenization significantly facilitates efficient learning of translating images into object and attribute descriptions compared to the traditional patch-level tokenization. Codes and models will be open-sourced at https://github.com/ChenDelong1999/subobjects.
Figure 1: The connection between tokenization methodologies in language modelling and image modelling. Our proposed subobject-level image tokenization is corresponded to the subword-level textual tokenization, which has been proven to be superior compared to other alternatives.
1 Introduction
In the field of Natural Language Processing (NLP), researchers have extensively investigated a variety of tokenization approaches over years Mielke2021words , but have converged to the subword-level tokenization recently, represented by Bytepair Encoding (BPE) Sennrich2016Neural , Unigram Kudo2018Subword , and the SentencePiece tokenizer Kudo2018SentencePiece adopted by many language models. Compared to the character-level and word-level counterparts, subword-level tokenization is better at capturing the morphology of a word which not only avoids the explosion of vocabulary size but also facilitates compositional generalization, e.g., the word “tokenization” and “generalization” can be decomposed into two tokens each, yielding “token” and “general” as the prefix, while sharing the same suffix token “-ization”. Additionally, subword-level tokenization brings out-of-vocabulary generalization, minimizing the production of unknown tokens during the tokenization process.
In Computer Vision (CV), since our visual world has not been discretized and optimized for communication as a language, it is more sparse, redundant, and noisy than textual data. Directly using raw pixels as input units (e.g., pixel-level tokenization adopted by ImageGPT Chen2020Generative ) leads to an excessively large number of tokens and unnecessary modeling of low-level relationships between neighboring pixels. The Vision Transformer (ViT) architecture Dosovitskiy2021Image , which is the dominant type of vision model in the current field, raises the tokenization level from pixels to square patches. However, as demonstrated in Fig. 1, such patch-level image tokenization corresponds to a character N-gram-level textual tokenization method, which tends to be both ineffective and inefficient due to the ignorance of semantic boundaries and humongous vocabulary size McNamee2004-hy ; Sennrich2016Neural . Similar to N-gram tokenization, the patch partitioning operation is not adaptive to the morphology of objects, ignoring the inherent pixel-grouping structure in the image.
In this paper, inspired by the performant subword-level text tokenization, we introduce the concept of “subobject”-level image tokenization, which lies in an intermediate level between objects and pixels, akin to subwords between words and characters. Subobjects are visual entities (e.g., parts of objects) with perceptually meaningful visual structures. It can be obtained through image segmentation, such as using Segment Anything Models (SAM) Kirillov2023Segment . The concept of subobject is related to the concept of superpixels Ren2003Learning ; Stutz2018Superpixels for low-level vision and the part segmentation He2022PartImageNet task in high-level vision, but it emphasizes the requirements of being semantically meaningful, open-vocabulary, and panoptic Kirillov2019Panoptic , also highlights the application to image tokenization and the connection to subword in NLP.
At the methodology level, this paper presents two types of neural architectures for effectively creating a learning system based on subobject-level image tokenization. Firstly, we propose a Sequence-to-sequence AutoEncoder (SeqAE) to compress subobject segments of varying sizes and shapes into compact embedding vectors. Compared to downsampling and fitting irregular segments into a square input window, SeqAE is able to reserve more information when handling segments with extreme aspect ratios. Secondly, we designed a simple yet effective architecture of Large Vision Language Model (LVLM), which incorporates subobject tokens into a Large Language Model (LLM) by treating them as textual subword tokens in new languages Wang2023Image .
Empirically, we trained a SeqAE on the SA-1B dataset Kirillov2023Segment , then trained an LVLM with subobject-level image tokenization based on the Phi-2 model Li2023Textbooks , using a synthetic captioning dataset created from CLEVR Johnson2017CLEVR . Our results demonstrate that subobject-level tokenization enables significantly accelerated vision-language learning compared to the standard ViT-style or Fuyu-style adeptFuyu8BMultimodal patch-level tokenization, while at the same time achieving higher accuracies in counting objects and recognizing visual attributes such as size, material, and shape by a large margin.
2 Method
In this section, we introduce our method of creating a learning system based on subobject-level tokenization. Our method consists of the following three steps: segmentation–obtaining subobject boundaries from images (section 2.1); embedding–converting raw pixels of subobject into compact vector embedding (section 2.2); modeling– building a model that takes embedded subobject tokens as inputs for vision-language learning (section 2.3).
Figure 2: The initial results from the “segment everything” mode of Segment Anything Model (SAM) Kirillov2023Segment leaves many pixels in blank. To avoid information loss, we perform post-processing of mask expansion and gap infilling to ensure the subobject segmentation covers all pixels.
2.1 Segment Everything into Subobjects
The desired subobject boundaries should be semantic meaningful, open-vocabulary, and comprehensive. Among other alternatives such as superpixel segmentation, semantic/instance/panoptic segmentation, the “segment everything” mode Zhang2023Faster ; Zhang2023MobileSAMv2 (also known as automated mask generation) of the Segment Anything Model (SAM) Kirillov2023Segment has more advantages on satisfying the requirements of being semantic meaningful and open-vocabulary. However, the results of “segment everything” are not guaranteed to be comprehensive. As shown in Fig. 2, there could be many pixels (shown in white) that are not covered by any mask in the “segment everything” results. These uncovered pixels usually correspond to the background or tiny gaps between neighboring segments. To ensure comprehensiveness, we conduct post-processing by applying convolution on the segmentation mask with a small kernel to expand the masks and filling the gaps (similar to binarizing blurred masks), then we do connected component labeling on pixels that are still not covered by any segments. Segments post-processed by such mask expansion and background filling can cover all pixels in the image and avoid any information loss.
2.2 SeqAE for Subobject Embedding
Subobject segments have irregular sizes and shapes. Although it’s possible to fit segments into square perception windows by padding to the longest side, it would be very inefficient when facing segments with extreme aspect ratios. Consider a Transformer encoder with a square perception window of 32$\times$32 pixels (i.e., 1024 pixels), it could only losslessly encode segments within the area on the right. However, as shown on the right, with same budget of 1024 context length, it is possible to encode wider and higher segments under the green curve (inverse proportion function) without any downsampling operations.
Figure 3: The architecture of our proposed Sequence-to-sequence AutoEncoder (SeqAE) for subobject embedding. SeqAE extracts compact latent variables from an image segment using learnable queries vectors and autoregressively decode the segment from the latent variables through the cross-attention mechanism.
We introduce Sequence-to-sequence AutoEncoder (SeqAE) to address this issue. In SeqAE, raw segment pixels and masks are flattened into data sequences to make full use of the context length. SeqAE is trained to compress subobject segments into compact embeddings via self-supervised autoencoding objective. As shown in Fig. 3, the encoder extracts a latent vector from the input data sequence, then the decoder reconstructs the inputs autoregressively. SeqAE shares many architectural similarities with the vanilla encoder-decoder Transformer language model for neural machine translation Vaswani2017Attention , but has the following two key modifications:
Real-valued regression instead of categorical prediction: In language models, the decoder predicts the categorical probabilistic distribution of the next token and compares it with ground truth discretized one-hot embedding via cross-entropy. For pixels, although it is possible to regard each RGB pixel as a $256^{3}$-way or three 256-way categorical distribution, it either results in an extremely large vocabulary size or increases the number of tokens by three times. More importantly, it ignores the inherent continual nature of pixel intensity and loses the relational information between pixel intensity values (e.g., the model does not know that red is closer to purple but much further to green). Therefore, in SeqAE we directly use the normalized real-valued pixel intensity as the data sequence, and then applies mean squared error (MSE) regression loss on the decoder output.
Extracting compact latents via learnable queries and bottleneck projector: Inspired by Perceiver Resampler Alayrac2022Flamingo , learnable query tokens are appended to the input data sequence. They interact with pixel tokens and integrate their information after going through the encoder layers. We further add a linear layer on top of these query tokens in the encoder’s last layer to reduce the dimension from $d_{\text{model}}\times n_{\text{query}}$ to $d_{\text{SeqAE}}$, and use the same layer (transposed) to reconstruct the query tokens, which the decoder can cross-attend to. This linear layer acts as a bottleneck to encourage information compression.
2.3 LVLM based on Subobject-level Image Tokenization
Figure 4: LLM to LVLM adaptation using our subobject-level image tokenization. We project the latent vector from SeqAE using a learnable linear weight and feed them into the LLM as parts of the input. We add a two-dimensional positional embedding to each subobject token to provide the position information of each subobject in the image.
Our methodology of inserting subobject tokens into LLMs is simple and straightforward. Inspired by Wang et al. Wang2023Image , we treat them as textual subword tokens in new languages, and then creating LVLM from LLM becomes equivalent to adding a foreign language to an LLM. As shown on the left side of Fig. 4, the process of obtaining Transformer’s input tokens from subword and subobject are conceptually similar. On the right side of Fig. 4, the subobject tokens are interleaved with subword tokens at the same level, only with a pair of new special tokens <SOI> and <EOI> marking the start and the end of subobject tokens from a single image.
However, considered as a foreign language, the image has one fundamental difference compared to natural languages, which is its dimensionality. To accommodate this unique nature, we make the following two technical modifications:
Additional positional embedding for subobject tokens: Since the original positional embedding existing in the LLM can only represent one-dimensional order relationships, we introduce additional two-dimensional positional embedding for subobject tokens. We use the absolute bounding box coordinates (in [x,y,w,h] format) of the segmentation mask to represent the position of subobjects. As shown in Fig. 4, we train a linear layer to project the bounding box into the same dimension of subobject tokens, then add them together and feed the result into Transformer layers.
No autoregressive prediction for subobject tokens: Images are two-dimensional projections of our three-dimensional visual world. These subobjects do not form any one-dimensional causal structure similar to natural language, making “next subobject token prediction” irrelevant. Therefore, during the LVLM training, we only calculate the cross-entropy loss on textual subword tokens while skipping the subobject tokens.
3 Experiment
Figure 5: Left: An example of our synthesized image caption from CLEVR Johnson2017CLEVR dataset. <image> indicates the position of subobject tokens. Middle: With subobject-level image tokenization, vision-language modelling becomes significantly faster. Right: The accuracy of generated descriptions in terms of size, material, shape, and count. Subobject-level method outperform standard patch-level baselin by a large margin.
3.1 Implementation of SeqAE
Model. Both the encoder and the decoder of SeqAE have 16 layers, where each layer is a standard Transformer layer with $d_{\text{model}}=768$, $d_{\text{FFN}}=4096$, and 12 attention heads. We use 16 learnable queries, each of them is a vector with a dimension of $d_{\text{model}}=768$. The encoder output $d_{\text{SeqAE}}$ is set to 768, and therefore the bottleneck projector is a linear layer of $16\times 768\rightarrow 768$. We use a context length of 1024 tokens, which can losslessly accommodate segments with width $\times$ heights $\leq 1024$ (e.g., 32$\times$32, 64$\times$16, 16$\times$64 segments, etc). Larger segments are down-sampled to 1024 tokens. The entire SeqAE model with the above configurations has a total of 327 million parameters.
Data. We use the SA-1B dataset Kirillov2023Segment to train the SeqAE model since it contains a larger-scale (1 billion) high-quality segmentation masks in of million images from various visual domains.
Training. We train SeqAE on the SA-1B dataset with a single-node 8$\times$×NVIDIA A100 (80GB) server. We use a batch size of 16 per GPU and a learning rate of 1e-5.
3.2 Implementation of LVLM
Model. We use the Phi-2 model [^1] https://huggingface.co/microsoft/phi-2, which is a base LLM with 2.7B parameters trained on high-quality textual data Li2023Textbooks . We apply LoRA Hu2022LoRA on its query/key/value projectors and FFN layers. Two trainable linear layers are newly initialized to project subobject embedding from the frozen SeqAE encoder and the bounding box coordinates to the input token space of the LLM.
Data. We create a synthetic image captioning dataset from CLEVR Johnson2017CLEVR , which contains scene graph annotations giving ground-truth locations, attributes, and relationships for objects. As shown in Fig. 5 left, we convert the scene graph annotations into textual descriptions of object counts, size, material, and shape. We arrange objects according to their positions (from left to right). We use 70k samples for training and evaluate 2k unseen samples.
Training. We use an effective batch size of 32 to train the model for 10 epochs. We use a cosine learning rate scheduler with a starting learning rate of 1e-4. MobileSAM-v2 Zhang2023MobileSAMv2 is used for subobject segmentation. Subobject segments and embeddings are cached in advance to boost training efficiency.
3.3 Results
We compare the LVLM based on the subobject-level image tokenizer to a patch-level baseline, which divides the input image into $32\times 32$ square patches – one of the most frequently used patch sizes in ViT models. Same as ViT, the patch token embedding is a linear layer that projects flattened patches ($32\times 32\times 3$) to the same dimension of SeqAE patents, and is also trained with autoencoding reconstruction objective. Other settings of this baseline are the same as the subobject-based LVLM.
The experimental results are presented in Fig. 5. On the left side, it shows that subobject-level tokenization enables a significantly faster decrease in training perplexity, showing that the same model can learn much faster when replacing patch-level tokenization with a subobject level one. On the right side, we present the evaluation of model-generated captions on 2k unseen testing images. We parse the generated captions (generated captions are 100% parseable), and respectively calculate the prediction accuracy of object size, material, shape, and counts with the ground truth. The results show that the model based on subobject-level tokenization outperforms the baseline by a large margin in all of the four aspects.
4 Conclusion
In this paper, we introduce subobject-level image tokenization which is a viable alternative to patch-level tokenization for vision-language learning. Our preliminary results demonstrate that compared to the standard patch-level baseline, subobject-level tokenization can accelerate vision-language learning, achieve higher accuracy in counting objects, and recognize visual attributes.

### v2

@@
 Inspired by the subword tokenization widely adopted in language models, we propose an image tokenizer at a subobject level, where the subobjects are represented by semantically meaningful image segments obtained by segmentation models (e.g., segment anything models).
-To implement a learning system based on subobject tokenization, we first introduced a Sequence-to-sequence AutoEncoder (SeqAE) to compress subobject segments of varying sizes and shapes into compact embedding vectors, then fed the subobject embeddings into a large language model for vision language learning.
+To implement a learning system based on subobject tokenization, we first introduced a Direct Segment Anything Model (DirectSAM) that efficiently produces comprehensive segmentation of subobjects, then embed subobjects into compact latent vectors and fed them into a large language model for vision language learning.
 Empirical results demonstrated that our subobject-level tokenization significantly facilitates efficient learning of translating images into object and attribute descriptions compared to the traditional patch-level tokenization.
-Codes and models will be open-sourced at https://github.com/ChenDelong1999/subobjects.
+Codes and models are open-sourced at https://github.com/ChenDelong1999/subobjects.
 Figure 1: The connection between tokenization methodologies in language modelling and image modelling.
@@
 1 Introduction
-In the field of Natural Language Processing (NLP), researchers have extensively investigated a variety of tokenization approaches over years Mielke2021words , but have converged to the subword-level tokenization recently, represented by Bytepair Encoding (BPE) Sennrich2016Neural , Unigram Kudo2018Subword , and the SentencePiece tokenizer Kudo2018SentencePiece adopted by many language models.
+In the field of Natural Language Processing (NLP), researchers have extensively investigated a variety of tokenization approaches over years [1], but have converged to the subword-level tokenization recently, represented by Bytepair Encoding (BPE) [2], Unigram [3], and the SentencePiece tokenizer [4] adopted by many language models.
 Compared to the character-level and word-level counterparts, subword-level tokenization is better at capturing the morphology of a word which not only avoids the explosion of vocabulary size but also facilitates compositional generalization, e.g., the word “tokenization” and “generalization” can be decomposed into two tokens each, yielding “token” and “general” as the prefix, while sharing the same suffix token “-ization”.
@@
 In Computer Vision (CV), since our visual world has not been discretized and optimized for communication as a language, it is more sparse, redundant, and noisy than textual data.
-Directly using raw pixels as input units (e.g., pixel-level tokenization adopted by ImageGPT Chen2020Generative ) leads to an excessively large number of tokens and unnecessary modeling of low-level relationships between neighboring pixels.
-The Vision Transformer (ViT) architecture Dosovitskiy2021Image , which is the dominant type of vision model in the current field, raises the tokenization level from pixels to square patches.
+Directly using raw pixels as input units (e.g., pixel-level tokenization adopted by ImageGPT [5]) leads to an excessively large number of tokens and unnecessary modeling of low-level relationships between neighboring pixels.
+The Vision Transformer (ViT) architecture [6], which is the dominant type of vision model in the current field, raises the tokenization level from pixels to square patches.
 However, as demonstrated in Fig.
-1, such patch-level image tokenization corresponds to a character N-gram-level textual tokenization method, which tends to be both ineffective and inefficient due to the ignorance of semantic boundaries and humongous vocabulary size McNamee2004-hy ; Sennrich2016Neural .
+1, such patch-level image tokenization corresponds to a character N-gram-level textual tokenization method, which tends to be both ineffective and inefficient due to the ignorance of semantic boundaries and humongous vocabulary size [7, 2].
 Similar to N-gram tokenization, the patch partitioning operation is not adaptive to the morphology of objects, ignoring the inherent pixel-grouping structure in the image.
@@
 Subobjects are visual entities (e.g., parts of objects) with perceptually meaningful visual structures.
-It can be obtained through image segmentation, such as using Segment Anything Models (SAM) Kirillov2023Segment .
-The concept of subobject is related to the concept of superpixels Ren2003Learning ; Stutz2018Superpixels for low-level vision and the part segmentation He2022PartImageNet task in high-level vision, but it emphasizes the requirements of being semantically meaningful, open-vocabulary, and panoptic Kirillov2019Panoptic , also highlights the application to image tokenization and the connection to subword in NLP.
+It can be obtained through image segmentation, such as using Segment Anything Models (SAM) [8].
+The concept of subobject is related to the concept of superpixels [9, 10] for low-level vision and the part segmentation [11] task in high-level vision, but it emphasizes the requirements of being semantically meaningful, open-vocabulary, and panoptic [12], also highlights the application to image tokenization and the connection to subword in NLP.
 At the methodology level, this paper presents two types of neural architectures for effectively creating a learning system based on subobject-level image tokenization.
-Firstly, we propose a Sequence-to-sequence AutoEncoder (SeqAE) to compress subobject segments of varying sizes and shapes into compact embedding vectors.
-Compared to downsampling and fitting irregular segments into a square input window, SeqAE is able to reserve more information when handling segments with extreme aspect ratios.
-Secondly, we designed a simple yet effective architecture of Large Vision Language Model (LVLM), which incorporates subobject tokens into a Large Language Model (LLM) by treating them as textual subword tokens in new languages Wang2023Image .
-Empirically, we trained a SeqAE on the SA-1B dataset Kirillov2023Segment , then trained an LVLM with subobject-level image tokenization based on the Phi-2 model Li2023Textbooks , using a synthetic captioning dataset created from CLEVR Johnson2017CLEVR .
-Our results demonstrate that subobject-level tokenization enables significantly accelerated vision-language learning compared to the standard ViT-style or Fuyu-style adeptFuyu8BMultimodal patch-level tokenization, while at the same time achieving higher accuracies in counting objects and recognizing visual attributes such as size, material, and shape by a large margin.
+Firstly, we propose a Direct Segment Anything Model (DirectSAM) that produces subobject boundaries at once instead of generating their masks sequentially.
+It significantly reduces the time cost for obtaining subobject segmentations, while improving the quality and comprehensiveness.
+Secondly, we designed a simple yet effective architecture of Large Vision Language Model (LVLM), which incorporates subobject tokens into a Large Language Model (LLM) by treating them as textual subword tokens in new languages [13].
+Empirically, we trained a DirectSAM on the SA-1B dataset [8], and trained an LVLM with subobject-level image tokenization based on the Phi-2 model [14], using a synthetic captioning dataset created from CLEVR [15].
+Our results demonstrate that subobject-level tokenization enables significantly accelerated vision-language learning compared to the standard ViT-style or Fuyu-style [16] patch-level tokenization, while at the same time achieving higher accuracies in counting objects and recognizing visual attributes such as size, material, and shape by a large margin.
 2 Method
@@
 Our method consists of the following three steps: segmentation–obtaining subobject boundaries from images (section 2.1); embedding–converting raw pixels of subobject into compact vector embedding (section 2.2); modeling– building a model that takes embedded subobject tokens as inputs for vision-language learning (section 2.3).
-Figure 2: The initial results from the “segment everything” mode of Segment Anything Model (SAM) Kirillov2023Segment leaves many pixels in blank.
-To avoid information loss, we perform post-processing of mask expansion and gap infilling to ensure the subobject segmentation covers all pixels.
-2.1 Segment Everything into Subobjects
+2.1 Subobject Segmentation with DirectSAM
 The desired subobject boundaries should be semantic meaningful, open-vocabulary, and comprehensive.
-Among other alternatives such as superpixel segmentation, semantic/instance/panoptic segmentation, the “segment everything” mode Zhang2023Faster ; Zhang2023MobileSAMv2 (also known as automated mask generation) of the Segment Anything Model (SAM) Kirillov2023Segment has more advantages on satisfying the requirements of being semantic meaningful and open-vocabulary.
-However, the results of “segment everything” are not guaranteed to be comprehensive.
-As shown in Fig.
-2, there could be many pixels (shown in white) that are not covered by any mask in the “segment everything” results.
-These uncovered pixels usually correspond to the background or tiny gaps between neighboring segments.
-To ensure comprehensiveness, we conduct post-processing by applying convolution on the segmentation mask with a small kernel to expand the masks and filling the gaps (similar to binarizing blurred masks), then we do connected component labeling on pixels that are still not covered by any segments.
-Segments post-processed by such mask expansion and background filling can cover all pixels in the image and avoid any information loss.
-2.2 SeqAE for Subobject Embedding
+Among other alternatives such as superpixel segmentation, semantic/instance/panoptic segmentation, the “segment everything” mode [17, 18] (also known as automated mask generation) of the Segment Anything Model (SAM) [8] has more advantages on satisfying the requirements of being semantic meaningful and open-vocabulary.
+However, the results of “segment everything” are not guaranteed to be comprehensive, there could be many pixels that are not covered by any segmentation mask.
+Also, running “segment everything” with SAM is time consuming, since it requires prompting SAM with a grid of points and ruining mask decoder for thousands of times.
+Attempts have been made for reducing the number of prompt [18], but it still needs hundreds times of decoder inference.
+The fundamental limitation of these prompting-based method is that their time complexity is $O(N)$ where $N$ is the number of subobject in an image–it’s impossible to guarantee efficiency and comprehensiveness at the same time.
+To address this limitation, we propose the Direct Segment Anything Model (DirectSAM), which reduces the time complexity from $O(N)$ to $O(1)$.
+DirectSAM directly predict the subobject boundary, thus it only requires a single pass of decoder inference.
+For training DirectSAM, the supervision signal can be converted from mask annotation via contour detection [19].
+Specifically, for each height×$\times$×width×$\times$×channel image, the target of DirectSAM prediction is a height×$\times$×width matrix, where all points corresponding to the extracted contours are one otherwise zero.
+Figure 2: Visualization of DirectSAM-based subobject segmentation.
+The black-white images are binarized boundary probabilities, while colored images fills each subobject area according to the input image.
+Figure 2 visualizes some subobject segmentations produced by DirectSAM, where test-time multi-scale augmentation is applied to improve the quality.
+The multi-scale augmentation, inspired by the recent Pyramid of Captions (PoCa) method [20], involves running DirectSAM on both global image and local patches and merging there predicted boundary probabilities.
+High-resolution version and more visualizations are available at our project website, and please see https://www.youtube.com/watch?v=tlNs7xUQ0x4 for a HD demo video.
+2.2 Subobject Embedding with SeqAE
 Subobject segments have irregular sizes and shapes.
@@
 However, as shown on the right, with same budget of 1024 context length, it is possible to encode wider and higher segments under the green curve (inverse proportion function) without any downsampling operations.
-Figure 3: The architecture of our proposed Sequence-to-sequence AutoEncoder (SeqAE) for subobject embedding.
-SeqAE extracts compact latent variables from an image segment using learnable queries vectors and autoregressively decode the segment from the latent variables through the cross-attention mechanism.
 We introduce Sequence-to-sequence AutoEncoder (SeqAE) to address this issue.
@@
 SeqAE is trained to compress subobject segments into compact embeddings via self-supervised autoencoding objective.
-As shown in Fig.
-3, the encoder extracts a latent vector from the input data sequence, then the decoder reconstructs the inputs autoregressively.
-SeqAE shares many architectural similarities with the vanilla encoder-decoder Transformer language model for neural machine translation Vaswani2017Attention , but has the following two key modifications:
+The encoder extracts a latent vector from the input data sequence, then the decoder reconstructs the inputs autoregressively.
+SeqAE shares many architectural similarities with the vanilla encoder-decoder Transformer language model for neural machine translation [21], but has the following two key modifications:
 Real-valued regression instead of categorical prediction: In language models, the decoder predicts the categorical probabilistic distribution of the next token and compares it with ground truth discretized one-hot embedding via cross-entropy.
@@
 Therefore, in SeqAE we directly use the normalized real-valued pixel intensity as the data sequence, and then applies mean squared error (MSE) regression loss on the decoder output.
-Extracting compact latents via learnable queries and bottleneck projector: Inspired by Perceiver Resampler Alayrac2022Flamingo , learnable query tokens are appended to the input data sequence.
+Extracting compact latents via learnable queries and bottleneck projector: Inspired by Perceiver Resampler [22], learnable query tokens are appended to the input data sequence.
 They interact with pixel tokens and integrate their information after going through the encoder layers.
@@
 Inspired by Wang et al.
-Wang2023Image , we treat them as textual subword tokens in new languages, and then creating LVLM from LLM becomes equivalent to adding a foreign language to an LLM.
+[13], we treat them as textual subword tokens in new languages, and then creating LVLM from LLM becomes equivalent to adding a foreign language to an LLM.
 As shown on the left side of Fig.
-4, the process of obtaining Transformer’s input tokens from subword and subobject are conceptually similar.
+3, the process of obtaining Transformer’s input tokens from subword and subobject are conceptually similar.
 On the right side of Fig.
-4, the subobject tokens are interleaved with subword tokens at the same level, only with a pair of new special tokens <SOI> and <EOI> marking the start and the end of subobject tokens from a single image.
+3, the subobject tokens are interleaved with subword tokens at the same level, only with a pair of new special tokens <SOI> and <EOI> marking the start and the end of subobject tokens from a single image.
 However, considered as a foreign language, the image has one fundamental difference compared to natural languages, which is its dimensionality.
@@
 As shown in Fig.
-4, we train a linear layer to project the bounding box into the same dimension of subobject tokens, then add them together and feed the result into Transformer layers.
+3, we train a linear layer to project the bounding box into the same dimension of subobject tokens, then add them together and feed the result into Transformer layers.
 No autoregressive prediction for subobject tokens: Images are two-dimensional projections of our three-dimensional visual world.
-These subobjects do not form any one-dimensional causal structure similar to natural language, making “next subobject token prediction” irrelevant.
+These subobjects do not form any one-dimensional causal structure similar to natural language, making “next subobject token prediction” counter-intuitive.
 Therefore, during the LVLM training, we only calculate the cross-entropy loss on textual subword tokens while skipping the subobject tokens.
 3 Experiment
-Figure 5: Left: An example of our synthesized image caption from CLEVR Johnson2017CLEVR dataset.
+Figure 4: Left: An example of our synthesized image caption from CLEVR [15] dataset.
 <image> indicates the position of subobject tokens.
@@
 Subobject-level method outperform standard patch-level baselin by a large margin.
+3.1 Implementation of DirectSAM
+Model.
+We use a Segformer [23] as the backbone of DirectSAM [^1] https://huggingface.co/nvidia/segformer-b5-finetuned-cityscapes-1024-1024, which has a total of 84.6M parameters.
+We replace the final multi-way classifier with a one-way classifier, and performs full-parameter fine-tuning.
+Data.
+We use the SA-1B dataset [8] to train the DirectSAM.
+The mask annotations are converted to boundaries via running opencv-implemented contour detection and plotting the extracted contours with a line width of 3.
+Random Gaussian blur is appllied with a probability of 0.25.
+Training.
+We train DirectSAM on the SA-1B dataset with a single-node 8$\times$×NVIDIA A100 (80GB) server.
+We first train it with a input resolution of 1024$\times$1024 for one epoch , then for another 0.6 epoch with 1800$\times$1800 resolution (the maximum resolution for data parallel training on 80GB GPUs).
+For the first 1024$\times$1024 epoch, we use a per GPU batch size of 4, gradient accumulation of steps of 4, and a learning rate of 4e-4.
+For the second 1800$\times$1800 epoch, we use a per GPU batch size of 1, gradient accumulation of steps of 8, and a learning rate of 2e-4.
+These two epochs respectively take around 15 days and 20 days.
 3.2 Implementation of SeqAE
@@
 The entire SeqAE model with the above configurations has a total of 327 million parameters.
-Data.
-We use the SA-1B dataset Kirillov2023Segment to train the SeqAE model since it contains a larger-scale (1 billion) high-quality segmentation masks in of million images from various visual domains.
+Data. We use the SA-1B dataset [8] to train the SeqAE model since it contains a larger-scale (1 billion) high-quality segmentation masks in of million images from various visual domains.
 Training. We train SeqAE on the SA-1B dataset with a single-node 8$\times$×NVIDIA A100 (80GB) server. We use a batch size of 16 per GPU and a learning rate of 1e-5.
@@
 Model.
-We use the Phi-2 model [^1] https://huggingface.co/microsoft/phi-2, which is a base LLM with 2.7B parameters trained on high-quality textual data Li2023Textbooks .
-We apply LoRA Hu2022LoRA on its query/key/value projectors and FFN layers.
+We use the Phi-2 model [^2] https://huggingface.co/microsoft/phi-2, which is a base LLM with 2.7B parameters trained on high-quality textual data [14].
+We apply LoRA [24] on its query/key/value projectors and FFN layers.
 Two trainable linear layers are newly initialized to project subobject embedding from the frozen SeqAE encoder and the bounding box coordinates to the input token space of the LLM.
 Data.
-We create a synthetic image captioning dataset from CLEVR Johnson2017CLEVR , which contains scene graph annotations giving ground-truth locations, attributes, and relationships for objects.
+We create a synthetic image captioning dataset from CLEVR [15], which contains scene graph annotations giving ground-truth locations, attributes, and relationships for objects.
 As shown in Fig.
@@
 We use a cosine learning rate scheduler with a starting learning rate of 1e-4.
-MobileSAM-v2 Zhang2023MobileSAMv2 is used for subobject segmentation.
+MobileSAM-v2 [18] is used for subobject segmentation [^3] Results of DirectSAM-based model will be presented in future updates..
 Subobject segments and embeddings are cached in advance to boost training efficiency.
@@
 The experimental results are presented in Fig.
-5.
+4.
 On the left side, it shows that subobject-level tokenization enables a significantly faster decrease in training perplexity, showing that the same model can learn much faster when replacing patch-level tokenization with a subobject level one.